OpenStudy (anonymous):

limit n tend to( n/2^n)the sequences is converge or diverg

OpenStudy (turingtest):

as n tends to infinity?

OpenStudy (turingtest):

use l'Hospiatals rule, do you know it?

OpenStudy (anonymous):

yes n tends to infinity

OpenStudy (turingtest):

I recommend using l'Hospital if you know it

OpenStudy (anonymous):

yes i know it hotial rule but procedure dont know . i mean step

OpenStudy (turingtest):

when you have a form like\[\pm\frac\infty\infty\]or\[\frac00\]you can take the derivative of the top and bottom, then try taking the limit again this is called l'Hospital's Rule what is the derivative of the numerator?

OpenStudy (anonymous):

top derivative is zero and bottom../?

OpenStudy (turingtest):

this is l'Hospital in action:\[\lim_{n\to\infty}\{a_n\}=\lim_{n\to\infty}{n\over2^n}=\lim_{n\to\infty}{\frac d{dn}(n)\over\frac d{dn}(2^n)}\]the derivative of n with respect to n is not zero please try again

OpenStudy (anonymous):

what is the derivative of 2^n

OpenStudy (turingtest):

\[\frac d{dx}(a^x)=a^x\ln a\]

OpenStudy (turingtest):

but you still haven't told me what the derivative of n is

OpenStudy (anonymous):

derivative of n is 1..?

OpenStudy (turingtest):

yes :)

OpenStudy (turingtest):

and what is the derivative of 2^n ?

OpenStudy (anonymous):

sorry sir i can't understand.. can u explain ..:(

OpenStudy (turingtest):

\[\frac d{dx}(a^x)=a^x\ln a\]you do not understand this formula?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

\[\frac d{dx}(a^x)=a^x\ln a\]examples:\[\frac d{dx}(5^x)=5^x\ln 5\]\[\frac d{dx}(132^x)=132^x\ln (132)\]\[\frac d{dx}(\pi^x)=\pi^x\ln\pi\]what is \(a\) in your case?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

right, so applying the formula\[\frac d{dx}(a^x)=a^x\ln a\]to\[\frac d{dx}(2^x)\]you get what?

OpenStudy (anonymous):

2^xln2

OpenStudy (turingtest):

exactly :)

OpenStudy (anonymous):

but in our case 2^nln(n) and the value of is infinity..

OpenStudy (turingtest):

why do you say ln(n) ? it will be ln2

OpenStudy (turingtest):

Remember how l'Hospital works: derivative of top and bottom separately when we have \(\pm\frac\infty\infty\) or \(\frac00\) so the derivative of the top is 1 the derivative of the bottom is (2^x)ln(2) so we have\[\lim_{n\to\infty}\{a_n\}=\lim_{n\to\infty}{n\over2^n}=\lim_{n\to\infty}{\frac d{dn}(n)\over\frac d{dn}(2^n)}=\lim_{n\to\infty}\frac1{2^n\ln 2}\]and that limit you can take, and it is...?

OpenStudy (anonymous):

2^n and n tends to infinity the 2^n i also infinity

OpenStudy (anonymous):

ln2 is 0.69

OpenStudy (turingtest):

ln(2) is a constant, so it doesn't tend to anything 2^n tends to infinity, so the whole fraction tends to...?

OpenStudy (anonymous):

infinity. bottom value is infinity ans1/infinity is zero. so our sequence is converg

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

than u so much what a explain.. thanks jazak allah

OpenStudy (turingtest):

you are welcome! please ask if you need more clarification

OpenStudy (anonymous):

nsin(pi/n)

OpenStudy (turingtest):

once again, I ask you to please post every question separately thank you

OpenStudy (turingtest):

it is also considered polite to click the button that says "best answer" if you like that answer not that I care, but it is how the site works

OpenStudy (anonymous):

thanks again i am very thankful to u